# ISAC Network Planning: Sensing Coverage Analysis and 3-D BS Deployment Optimization

### v1

@@ Abstract
Integrated sensing and communication (ISAC) networks strive to deliver both high-precision target localization and high-throughput data services across the entire coverage area. In this work, we examine the fundamental trade-off between sensing and communication from the perspective of base station (BS) deployment. Furthermore, we conceive a design that simultaneously maximizes the target localization coverage, while guaranteeing the desired communication performance. In contrast to existing schemes optimized for a single target, an effective network-level approach has to ensure consistent localization accuracy throughout the entire service area. While employing time-of-flight (ToF) based localization, we first analyze the deployment problem from a localization-performance coverage perspective, aiming for minimizing the area Cramér-Rao Lower Bound (A-CRLB) to ensure uniformly high positioning accuracy across the service area. We prove that for a fixed number of BSs, uniformly scaling the service area by a factor $\kappa$ increases the optimal A-CRLB in proportion to $\kappa^{2\beta}$, where $\beta$ is the BS-to-target pathloss exponent. Based on this, we derive an approximate scaling law that links the achievable A-CRLB across the area of interest to the dimensionality of the sensing area. We also show that cooperative BSs extends the coverage but yields marginal A-CRLB improvement as the dimensionality of the sensing area grows. By exploiting the invariance properties discovered with respect to the displacement, rotation, and symmetric projection deformation, we derive a deployment-invariant structure for conceiving a low-complexity framework for ISAC network deployment. We then formulate the joint sensing-communication optimization problem and present a Majorization-Minimization algorithm for designing high-quality deployment solutions. Extensive simulations demonstrate that our framework significantly enhances sensing coverage, while maintaining the desired communication throughput.
Index Terms:
Integrated sensing and communication, multi-cell networks, network performance analysis, stochastic geometry, antenna allocation, cooperative sensing and communication.
@@ I Introduction
Recent studies have begun exploring the new degrees of freedom (DoF) in managing the trade-offs between communication and sensing at the network level [18, 17, 19, 20, 21]. In particular, the authors of [15] and [22] present a systematic study of the emerging network-level performance metrics and DoF, laying the foundation for understanding and optimizing cooperative ISAC systems. As an example, [18] investigates how to optimize the number of cooperating BSs to maximize the attainable networked sensing and networked communication performance under specific backhaul constraints. In addition, [20] proposed a system-level beam alignment scheme that leverages synchronization signal block and time-frequency pattern design for dramatically reducing the beam misalignment probability in THz/mmWave networks. Most recently, [21] investigated antenna topology optimization in ISAC networks with randomly distributed targets, users, and BSs, introducing a new DoF in the S&C tradeoff and showing that the optimal topology may lie between centralized massive MIMO and distributed cell free topologies, depending on the localization method and path loss exponent. Overall, the above works mainly focus on waveform design and resource allocation, while the problem of optimizing the network deployment to enhance cooperation efficiency remains largely unexplored.
Traditional cellular communication networks have historically been deployed to maximize communication performance, optimizing coverage probability and data throughput [23]. However, as future networks shift toward integrated functionalities, the dual demands of reliable data transmission and high-accuracy environmental or target sensing become increasingly critical [24]. While communication systems are typically interference limited, the opportunities for multi-static sensing mean that sensing systems are not interference limited, which creates a fundamental paradigm shift for the ISAC network deployment. On the other hand, inadequate BS placement can induce severe sensing blind spots, undermining seamless localization in mission-critical applications, such as autonomous driving and dynamic target tracking in smart cities. To date, no systematic study has addressed how to (i) reconfigure existing BS layouts or (ii) augment them with additional sites for jointly optimizing sensing fidelity and communication quality. Filling this knowledge gap is essential to fully exploit the network-level DoF identified in prior research [18, 17, 19, 20], ensuring that future deployments meet the dual imperatives of robust data transmission and precise environmental awareness.
@@
Extending deployment strategies from isolated, single-target sensor placements to full area-wide S&C coverage requires incorporating spatially varying path loss and geometric gain for every target location, which directly extends and reshapes the design space. Moreover, evaluating sensing performance via metrics such as the CRLB introduces intricate couplings among BS positions and network parameters, rendering analytical tractability elusive and necessitating a complete redesign of optimization algorithms [29]. The resultant interdependence of BS locations not only amplifies the computational burden of determining an optimal configuration but also invalidates conventional distance-based cooperation criteria. In cooperative multi-BS localization, the nuanced interactions among BSs directly influence both communication reliability and localization precision. Importantly, the scaling law that links the number of BSs to localization accuracy across an entire area of interest remains unexplored, because existing studies have been confined to the single-target scenario. Consequently, achieving robust and network-wide ISAC performance requires novel deployment and cooperation frameworks that explicitly account for spatial diversity, parameter coupling, and the cooperative dynamics of modern localization techniques.
@@
To overcome these challenges, we formulate the joint deployment problem as the twinned optimization of localization and communication coverage, aiming for minimizing the target localization accuracy across the entire target area, while simultaneously ensuring the required communication quality. By examining the structure of the CRLB, we unveil that the optimal placement exhibits displacement, rotation, and symmetric projection invariance that we detail in Section III. These properties allow us to collapse the high-dimensional search space into a much smaller, representative subset. Such invariance-driven reduction not only yields rigorous performance bounds, characterizing, e.g., the minimum achievable average CRLB as a function of network size and area geometry, but also admits low-complexity algorithms for practical ISAC network deployments. This framework directly addresses the trade-off between communication and sensing by decoupling the interdependencies among BSs, paving the way for scalable multi-cell ISAC deployments. Then, based on the Majorization-Minimization (MM) optimization framework, we transform the problem formulated to sequential sub-problems and effectively decouple the complex interrelations among BS positions. This decoupling not only simplifies the underlying optimization problem but also allows us to have more efficient algorithmic solutions. The main contributions of this paper are summarized as follows:
@@
  • We propose a cooperative ISAC network architecture that tightly integrates multi-static radar sensing with coordinated multi-point (CoMP) data transmission, by optimizing BS deployment to reveal the fundamental trade-offs between sensing and communication performance. Beyond traditional distributed MIMO radar deployments for individual target, our design simultaneously guarantees area-wide high-throughput communications and high-precision sensing.
@@
  • Employing time-of-flight (ToF) localization, we first establish that the optimal BS deployment strategy for sensing performance enhancement over an area admits three provable invariances, displacement, rotation, and reflection symmetry, thereby showing that any Euclidean transformation of a candidate layout preserves its sensing–communication performance. Moreover, we prove that for a fixed number of BSs, uniformly scaling the service area by a factor $\kappa$ increases the optimal area CRLB (A-CRLB) in proportion to $\kappa^{2\beta}$, where $\beta$ denotes the pathloss exponent between BSs and targets. Leveraging the conclusions derived, we can find an optimal solution for a new area based on the foundational optimal BS deployment strategy through displacement, rotations, symmetric projection, and area scaling.
[…]
@@
  • We reformulate the general BS deployment optimization for joint sensing–communication design within the MM framework, converting it into a sequence of tractable subproblems. By embedding a trust-region solver in each MM iteration, the resultant algorithm reliably converges to high-quality stationary solutions, while maintaining only modest per-iteration computational complexity. Extensive simulations demonstrate that the proposed deployment strategy is capable of enhancing the sensing-coverage probability by at least 50%, while maintaining the desired communication throughput.
@@
Notation: Lower-case letters in bold font will denote deterministic vectors. For instance, $X$ and ${\bf{X}}$ denote a one-dimensional (scalar) random variable and a random vector (containing more than one element), respectively. Similarly, $x$ and ${\bf{x}}$ denote scalar and deterministic vectorial values, respectively. ${\rm{E}}_{x}[\cdot]$ represents statistical expectation over the distribution of $x$, and []delimited-[][$\cdot$][ ⋅ ] represents a variable set. Let ${\bf{1}}^{T}=[1,1,\dots,1]\in\mathbb{R}^{1\times N}$ and ${\bf{e}}_{1}^{T}=[1,0,\dots,0]\in\mathbb{R}^{1\times N}$.
@@ II-A Network Model
As shown in Fig. 1, BSs within a given area form a cooperative cluster for both communication and sensing. For communication, these BSs can adopt non-coherent joint CoMP transmission, transmitting identical data streams without requiring strict phase synchronization, thereby enhancing received power, while maintaining low coordination overhead [18]. For sensing, the BSs collaborate as a distributed multi-static MIMO radar system, using code-division multiplexing (CDM) to ensure orthogonality among transmitted waveforms and achieve accurate localization under non-coherent signal processing. [^1] By avoiding phase-level synchronization, our cooperative ISAC strategy strikes a practical balance between system complexity and the performance of ISAC networks.
@@
d-dimensional area CRLB Optimal deployment $\frac{1}{N^{2\beta/d}}$star Proposed
@@
In this study, we explore the optimal BS deployment strategy for cooperative ISAC networks, to reveal that the deployment of BSs brings a new DoF to balance the sensing and communication performance. In this cooperative service area, each BS designs the transmit precoding for sending the information signal $s^{c}$ to the single-antenna communication user served, together with a dedicated radar signal $s^{s}_{n}$ for the detected target. Then, the communication user and the sensing target can be collaboratively served by $N$ BSs, indexed by $n$, and each BS has $M_{\rm{t}}$ transmit antennas and $M_{\rm{r}}$ receive antennas. Let us assume that the transmitted radar signals $\{s_{n}^{s}\}_{n=1}^{N}$ of the BSs in the cooperative sensing cluster are approximately orthogonal for any time delay of interest by applying CDM only to the sensing signal $s^{s}_{n}$. Accordingly, we spread it as $\tilde{s}^{s}_{n}=\mathbf{c}_{n}\,s^{s}_{n}$, where $\|\mathbf{c}_{n}\|^{2}=1,\quad\mathbb{E}[\mathbf{c}_{n}\mathbf{c}_{m}^{H}]=\delta_{nm}$. For notational simplicity, we continue to use $s^{s}_{n}$ to represent the spread signal $\mathbf{c}_{n}s^{s}_{n}$ in the remainder of the discussion. It is assumed that $\mathrm{E}[s^{s}_{n}(s^{c}_{n})^{H}]=0$, which is consistent with the assumptions in [30, 31, 32]. Upon letting ${\mathbf{s}_{n}=\left[s^{s}_{n},s^{c}_{n}\right]^{T}}$, we have $\mathrm{E}\left[\mathbf{s}_{n}\mathbf{s}_{n}^{H}\right]=\mathbf{I}_{2}$. To facilitate our analysis, when a BS is connected to multiple users and targets, it is assumed that they are scheduled to different orthogonal time/frequency resource blocks. This ensures that each BS serves only one user and one target in a given time/frequency slot. [^2] Nevertheless, our analytical framework based on this assumption may be readily extended to multi-users and multi-targets association to each time/frequency resource blocks unless they are closely located within serving area. Then, the signal transmitted by the $n$th BS is given by
${\mathbf{x}}_{n}=\mathbf{W}_{n}\mathbf{s}_{n}={\bf{w}}^{c}_{n}s^{c}_{n}+{\bf{w}}^{s}_{n}s^{s}_{n},\vspace{-1.5mm}$ (1)
@@
where ${\bf{w}}^{c}_{n}$ and ${\bf{w}}^{s}_{n}\in{\mathbb{C}}^{M_{\mathrm{t}}\times 1}$ are normalized transmit beamforming vectors, i.e., $\|{\bf{w}}^{c}_{n}\|^{2}=p^{c}$ and $\|{\bf{w}}^{s}_{n}\|^{2}=p^{s}$. Here, $p^{s}$ and $p^{c}$ respectively represent the transmit power of the sensing and communication signals, and $\mathbf{W}_{n}=\left[\mathbf{w}^{c}_{n},\mathbf{w}^{s}_{n}\right]\in{\mathbb{C}}^{M_{\mathrm{t}}\times 2}$ is the transmit precoding matrix of BS $n$. The location of BS $n$ can be represented by ${\bf{b}}_{n}=[x_{n}^{b},y_{n}^{b},z_{n}^{b}]^{T}\in\mathbb{R}^{3},n=1,2,\dots,N$. To eliminate intra-BS interference from sensing to communications and simplify our analysis, zero-forcing (ZF) beamforming is employed. The beamforming matrix is constructed as
${\bf{W}}_{n}={\tilde{\mathbf{W}}}_{n}\left(\sqrt{\operatorname{diag}\left({\tilde{\mathbf{W}}}_{n}^{H}{\tilde{\mathbf{W}}}_{n}\right)}\right)^{-1},\vspace{-1.5mm}$ (2)
@@
where ${\tilde{\mathbf{W}}}_{n}={{\bf{H}}^{H}_{n}}{\left({\bf{H}}_{n}{\bf{H}}_{n}^{H}\right)^{-1}}$ and $\mathbf{H}_{n}=\begin{bmatrix}\mathbf{h}_{n,c}^{H}\\[2.15277pt] \mathbf{a}^{H}(\theta_{n})\end{bmatrix}\;\in\;\mathbb{C}^{2\times M_{t}}$. Here, $\mathbf{h}_{n,c}^{H}\in\mathbb{C}^{1\times M_{\rm{t}}}$ represents the communication channel spanning from BS $n$ to the served user, and $\mathbf{a}^{H}(\Omega_{n})$ corresponds to the sensing channel impinging from BS $n$ to the target. Here, $\Omega_{n}$ denotes the spatial frequency associated with angles of arrival from BS $n$ to the target point ${\bf{t}}=[x^{t},y^{t},z^{t}]^{T}\in\mathbb{R}^{3}$ along the antenna array. With the aid of ZF beamforming, intra-cell communication interference is minimized since all cooperating BSs serve the same communication user, while their sensing beams avoid that user.

### v2

@@ Abstract
Integrated sensing and communication (ISAC) networks strive to deliver both high‐precision target localization and high‐throughput data services across the entire coverage area. In this work, we examine the fundamental trade-off between sensing and communication from the perspective of base station (BS) deployment. Furthermore, we conceive a design that simultaneously maximizes the target localization coverage, while guaranteeing the desired communication performance. In contrast to existing schemes optimized for a single target, an effective network-level approach has to ensure consistent localization accuracy throughout the entire service area. While employing time-of-flight (ToF) based localization, we first analyze the deployment problem from a localization-performance coverage perspective, aiming for minimizing the area Cramér-Rao Lower Bound (A-CRLB) to ensure uniformly high positioning accuracy across the service area. We prove that for a fixed number of BSs, uniformly scaling the service area by a factor $\kappa$ increases the optimal A-CRLB in proportion to $\kappa^{2\beta}$, where $\beta$ is the BS‐to‐target pathloss exponent. Based on this, we derive an approximate scaling law that links the achievable A-CRLB across the area of interest to the dimensionality of the sensing area. We also show that cooperative BSs extend the coverage but yield marginal A-CRLB improvement as the dimensionality of the sensing area grows. By exploiting the invariance properties discovered with respect to the displacement, rotation, and symmetric projection deformation, we derive a deployment-invariant structure for conceiving a low-complexity framework for ISAC network deployment. We then formulate the joint sensing-communication optimization problem and present a Majorization-Minimization algorithm for designing high-quality deployment solutions. Extensive simulations demonstrate that our framework significantly enhances sensing coverage, while maintaining the desired communication throughput.
@@ I Introduction
Recent studies have begun exploring the new degrees of freedom (DoF) in managing the trade-offs between communication and sensing at the network level [18, 17, 19, 20, 21]. In particular, the authors of [15] and [22] present a systematic study of the emerging network-level performance metrics and DoF, laying the foundation for understanding and optimizing cooperative ISAC systems. As an example, [18] investigates how to optimize the number of cooperating BSs to maximize the attainable networked sensing and networked communication performance under specific backhaul constraints. Coordinated precoding can repurpose inter-cell interference as useful spatial degrees of freedom, substantially improving sensing performance. In addition, [20] proposed a system-level beam alignment scheme that leverages synchronization signal block and time-frequency pattern design for dramatically reducing the beam misalignment probability in THz/mmWave networks. Most recently, [21] investigated antenna topology optimization in ISAC networks with randomly distributed targets, users, and BSs, introducing a new DoF in the S&C tradeoff and showing that the optimal topology may lie between centralized massive MIMO and distributed cell free topologies, depending on the localization method and path loss exponent. Overall, the above works mainly focus on waveform design and resource allocation, while the problem of optimizing the network deployment to enhance cooperation efficiency remains largely unexplored.
Traditional cellular communication networks have historically been deployed to maximize communication performance, optimizing coverage probability and data throughput [23]. However, as future networks shift toward integrated functionalities, the dual demands of reliable data transmission and high-accuracy environmental or target sensing become increasingly critical [24]. While communication systems are typically interference-limited, the opportunities for multi-static sensing mean that sensing systems are not interference limited, which creates a fundamental paradigm shift for the ISAC network deployment. On the other hand, inadequate BS placement can induce severe sensing blind spots, undermining seamless localization in mission-critical applications, such as autonomous driving and dynamic target tracking in smart cities. To date, no systematic study has addressed how to (i) reconfigure existing BS layouts or (ii) augment them with additional sites for jointly optimizing sensing fidelity and communication quality. Filling this knowledge gap is essential to fully exploit the network-level DoF identified in prior research [18, 17, 19, 20], ensuring that future deployments meet the dual imperatives of robust data transmission and precise environmental awareness.
@@
Extending deployment strategies from isolated, single-target sensor placements to full area-wide S&C coverage requires incorporating spatially varying path loss and geometric gain for every target location, which directly extends and reshapes the design space. Moreover, evaluating sensing performance via metrics such as the CRLB introduces intricate couplings among BS positions and network parameters, rendering analytical tractability elusive and necessitating a complete redesign of optimization algorithms [29]. The resultant interdependence of BS locations not only amplifies the computational burden of determining an optimal configuration but also invalidates conventional distance‐based cooperation criteria. In cooperative multi‐BS localization, the nuanced interactions among BSs directly influence both communication reliability and localization precision. Importantly, the scaling law that links the number of BSs to localization accuracy across an entire area of interest remains unexplored, because existing studies have been confined to the single-target scenario. Consequently, achieving robust and network‐wide ISAC performance requires novel deployment and cooperation frameworks that explicitly account for spatial diversity, parameter coupling, and the cooperative dynamics of modern localization techniques.
@@
To overcome these challenges, we formulate the joint deployment problem as the twinned optimization of localization and communication coverage, aiming for minimizing the target localization error across the entire target area, while simultaneously ensuring the required communication quality. By examining the structure of the CRLB, we unveil that the optimal placement exhibits displacement, rotation, and symmetric projection invariance that we detail in Section III. These properties allow us to collapse the high-dimensional search space into a much smaller, representative subset. Such invariance-driven reduction not only yields rigorous performance bounds, characterizing, e.g., the minimum achievable average CRLB as a function of network size and area geometry, but also admits low-complexity algorithms for practical ISAC network deployments. This framework directly addresses the trade-off between communication and sensing by decoupling the interdependencies among BSs, paving the way for scalable multi-cell ISAC deployments. Then, based on the Majorization-Minimization (MM) optimization framework, we transform the problem formulated to sequential sub-problems and effectively decouple the complex interrelations among BS positions. This decoupling not only simplifies the underlying optimization problem but also allows us to have more efficient algorithmic solutions. The main contributions of this paper are summarized as follows:
@@
  • We propose a cooperative ISAC network architecture that tightly integrates multi-static radar sensing with coordinated multi-point (CoMP) data transmission, by optimizing BS deployment to reveal the fundamental trade-offs between sensing and communication performance. Beyond traditional distributed MIMO radar deployments for individual target, our design simultaneously guarantees area‐wide high‐throughput communications and high-precision sensing.
@@
  • Employing time-of-flight (ToF) localization, we first establish that the optimal BS deployment strategy for sensing performance enhancement over an area admits three provable invariances, displacement, rotation, and reflection symmetry, thereby showing that any Euclidean transformation of a candidate layout preserves its sensing-communication performance. Moreover, we prove that for a fixed number of BSs, uniformly scaling the service area by a factor $\kappa$ increases the optimal area CRLB (A-CRLB) in proportion to $\kappa^{2\beta}$, where $\beta$ denotes the pathloss exponent between BSs and targets. Leveraging the conclusions derived, we can find an optimal solution for a new area based on the foundational optimal BS deployment strategy through displacement, rotations, symmetric projection, and area scaling.
  • We derive an asymptotic scaling law for the minimum A-CRLB $\frac{1}{N^{\,2\beta/d}}$, where $N$ denotes the number of cooperating ISAC nodes and $d\in\{1,2,3\}$ represents the spatial dimensionality of the sensing area. This law implies that, for a given increase in $N$, the reduction in sensing error decays more slowly as $d$ grows. Consequently, the incremental gain in coverage afforded by each additional node diminishes in higher-dimensional monitoring scenarios, underscoring an inherent trade-off between achievable sensing accuracy and the spatial extent of the surveillance area in the design of practical ISAC networks.
@@
  • We reformulate the general BS deployment optimization for joint sensing-communication design within the MM framework, converting it into a sequence of tractable subproblems. By embedding a trust‐region solver in each MM iteration, the resultant algorithm reliably converges to high‐quality stationary solutions, while maintaining only modest per‐iteration computational complexity. Extensive simulations demonstrate that the proposed deployment strategy is capable of enhancing the sensing‐coverage probability by at least 50%, while maintaining the desired communication throughput.
@@
Notation: Lower-case letters in bold font will denote deterministic vectors. For instance, $X$ and $\bf{X}$} denote a one-dimensional (scalar) random variable and a random vector (containing more than one element), respectively. Similarly, $x$ and $\bf{x}$} denote scalar and deterministic vectorial values, respectively. ${\rm{E}}_{x}[\cdot]$ represents statistical expectation over the distribution of $x$, and $\cdot$] represents a variable set. Let ${\bf{1}}^{T}\triangleq[1,1,\dots,1]\in\mathbb{R}^{1\times N}$ and $\mathbf{e}_{n}^{T}\triangleq[\,0,\ldots,0,\underbrace{1}_{n\text{th}},0,\ldots,0\,]\in\mathbb{R}^{1\times N}$.
@@ II-A Network Model
As shown in Fig. 1, BSs within a given area form a cooperative cluster for both communication and sensing. For communication, these BSs can adopt non-coherent joint CoMP transmission, transmitting identical data streams without requiring strict phase synchronization, thereby enhancing received power, while maintaining low coordination overhead [18]. For sensing, the BSs collaborate as a distributed multi-static MIMO radar system, using code-division multiplexing (CDM) to ensure orthogonality among transmitted waveforms and achieve accurate localization under non-coherent signal processing [30, 31]. [^1] By avoiding phase-level synchronization, our cooperative ISAC strategy strikes a practical balance between system complexity and the performance of ISAC networks.
@@
d-dimensional area CRLB Optimal deployment $\frac{1}{N^{2\beta/d}}$ $\star$ Proposed
@@
In this study, we explore the optimal BS deployment strategy for cooperative ISAC networks, to reveal that the deployment of BSs brings a new DoF to balance the sensing and communication performance. In this cooperative service area, each BS designs the transmit precoding for sending the information signal $s^{c}$ to the single-antenna communication user served, together with a dedicated radar signal $s^{s}_{n}$ for the detected target. Then, the communication user and the sensing target can be collaboratively served by $N$ BSs, indexed by $n$, and each BS has $M_{\rm{t}}$ transmit antennas and $M_{\rm{r}}$ receive antennas. Let us assume that the transmitted radar signals $\{s_{n}^{s}\}_{n=1}^{N}$ of the BSs in the cooperative sensing cluster are approximately orthogonal for any time delay of interest by applying CDM only to the sensing signal $s^{s}_{n}$. Accordingly, we spread it as $\tilde{s}^{s}_{n}=\mathbf{c}_{n}\,s^{s}_{n}$, where $\mathbf{c}_{n}^{H}\mathbf{c}_{m}=\delta_{nm}$, $\|\mathbf{c}_{n}\|^{2}=1$.
For notational simplicity, we continue to use $s^{s}_{n}$ to represent the spread signal $\mathbf{c}_{n}s^{s}_{n}$ in the remainder of the discussion. It is assumed that $\mathrm{E}[s^{s}_{n}(s^{c}_{n})^{H}]=0$, which is consistent with the assumptions in [32, 33, 34]. Upon letting $\mathbf{s}_{n}=\left[s^{s}_{n},s^{c}_{n}\right]^{T}$}, we have $\mathrm{E}\left[\mathbf{s}_{n}\mathbf{s}_{n}^{H}\right]=\mathbf{I}_{2}$. To facilitate our analysis, when a BS is connected to multiple users and targets, we assign them to orthogonal time or frequency resource blocks so that, within a certain resource block, each BS serves only one user and one target. For the block designated to a specific target or region of interest, all BSs are scheduled to serve that target simultaneously, enabling cooperative multistatic sensing. [^2] Nevertheless, our analytical framework based on this assumption may be readily extended to multi-users and multi-targets association to each time/frequency resource blocks unless they are closely located within serving area. Then, the signal transmitted by the $n$th BS is given by
${\mathbf{x}}_{n}=\mathbf{W}_{n}\mathbf{s}_{n}={\bf{w}}^{c}_{n}s^{c}_{n}+{\bf{w}}^{s}_{n}s^{s}_{n},\vskip 0.0pt$ (1)
@@
where ${\bf{w}}^{c}_{n}$ and ${\bf{w}}^{s}_{n}\in{\mathbb{C}}^{M_{\mathrm{t}}\times 1}$ are beamforming vectors, with $\|{\bf{w}}^{c}_{n}\|^{2}=p^{c}$ and $\|{\bf{w}}^{s}_{n}\|^{2}=p^{s}$. Here, $p^{s}$ and $p^{c}$ respectively represent the transmit power of the sensing and communication signals, and $\mathbf{W}_{n}=\left[\mathbf{w}^{c}_{n},\mathbf{w}^{s}_{n}\right]\in{\mathbb{C}}^{M_{\mathrm{t}}\times 2}$ is the transmit precoding matrix of BS $n$. The location of BS $n$ can be represented by ${\bf{b}}_{n}=[x_{n}^{b},y_{n}^{b},z_{n}^{b}]^{T}\in\mathbb{R}^{3},n=1,2,\dots,N$. To eliminate intra-BS interference from sensing to communications and simplify our analysis, zero-forcing (ZF) beamforming is employed. The beamforming matrix is constructed as
${\bf W}_{n}=\tilde{\bf W}_{n}\left(\sqrt{\mathrm{diag}\left(\tilde{\bf W}_{n}^{H}\tilde{\bf W}_{n}\right)}\right)^{-1}\mathrm{diag}\big(\sqrt{p^{c}},\sqrt{p^{s}}\big).$ (2)
@@
where ${\tilde{\mathbf{W}}}_{n}={{\bf{H}}^{H}_{n}}{\left({\bf{H}}_{n}{\bf{H}}_{n}^{H}\right)^{-1}}$ and $\mathbf{H}_{n}=\begin{bmatrix}\mathbf{h}_{n,c}^{H}\\[2.15277pt] \mathbf{a}^{H}_{M_{\mathrm{t}}}(\Omega_{n})\end{bmatrix}\;\in\;\mathbb{C}^{2\times M_{t}}$. Here, $\mathbf{h}_{n,c}^{H}\in\mathbb{C}^{1\times M_{\rm{t}}}$ represents the communication channel spanning from BS $n$ to the served user, and $\mathbf{a}^{H}(\Omega_{n})$ corresponds to the sensing channel impinging from BS $n$ to the target. Here, $\Omega_{n}$ denotes the spatial frequency associated with angles of arrival from BS $n$ to the target point ${\bf{t}}=[x^{t},y^{t},z^{t}]^{T}\in\mathbb{R}^{3}$ along the antenna array, for an uniform linear array aligned with the $x$-axis with spacing $\lambda/2$. With azimuth-elevation angles $(\varphi_{n},\theta_{n})$ from BS $n$ to the target, we have $\Omega_{n}=\pi\,\sin\theta_{n}\cos\varphi_{n}$. The spatial frequency $\Omega_{n}$ is the per-element phase increment induced by the direction $(\varphi_{n},\theta_{n})$ projected onto the array axis, i.e., $\mathbf{a}_{M_{\mathrm{t}}}(\Omega_{n})=\big[1,\ e^{\mathrm{j}\Omega_{n}},\ \ldots,\ e^{\mathrm{j}(M_{\mathrm{t}}-1)\Omega_{n}}\big]^{\top}$. With the aid of ZF beamforming, intra-cell communication interference is minimized since all cooperating BSs serve the same communication user, while their sensing beams avoid that user.